# Periodic Korteweg-de Vries soliton potentials generate magnetic field strength with excellent quasisymmetry

### v2

@@ Abstract
Quasisymmetry (QS) is a hidden symmetry of the magnetic field strength, $|\mathbf{B}|$, that confines charged particles effectively in a three-dimensional toroidal plasma equilibrium. Here, we show that QS has a deep connection to the underlying symmetry that makes solitons possible. Our approach uncovers a hidden lower dimensionality of $|\mathbf{B}|$ on a magnetic flux surface, which could make stellarator optimization schemes significantly more efficient. Recent numerical breakthroughs (M. Landreman and E. Paul, Phys. Rev. Lett. 128, 035001 (2022)) have yielded configurations with excellent volumetric QS and surprisingly low magnetic shear. Our approach elucidates why the magnetic shear is low in these configurations. Furthermore, we deduce an upper bound on the maximum toroidal volume that can be quasisymmetric and verify it for the Landreman-Paul precise quasiaxisymmetric (QA) stellarator configuration. In the neighborhood of the outermost surface, we show that the $|\mathbf{B}|$ approaches the form of the 1-soliton reflectionless potential (I. Gjaja and A. Bhattacharjee, Phys. Rev. Lett. 68, 2413 (1992)). We present three independent approaches to demonstrate that quasisymmetric $|\mathbf{B}|$ is described by well-known integrable systems such as the Korteweg-de Vries (KdV) equation. The first approach is weakly nonlinear multiscale perturbation theory, which highlights the crucial role that magnetic shear plays in QS. We show that the overdetermined problem of finding quasisymmetric vacuum fields admits solutions for which the rotational transform is not free but highly constrained. We obtain the KdV equation (and, more specifically, Gardner’s equation for certain choices of parameters). Our second approach is non-perturbative and based on ensuring single-valuedness of $|\bm{B}|$, which directly leads to its Painlevé property shared by the KdV equation. Our third approach uses machine learning, trained on a large dataset of numerically optimized quasisymmetric stellarators. We robustly recover the KdV (and Gardner’s) equation from the data.
@@ I Introduction
Physical systems possessing a hidden symmetry attract serious attention [1, 2, 3, 4, 5, 6] because of the far-reaching impact of the symmetry on the physical properties of the systems. The hidden symmetry manifests only in special coordinates, which are constructed, along with the invariants, as part of the solution to the problem. In the presence of hidden symmetries, the partial differential equations (PDEs) governing classical dynamical fields often reduce to integrable PDEs [5, 7, 6]. An example of the latter is the well-known Korteweg-de Vries (KdV) equation with soliton solutions [8, 9].
In plasma physics, a hidden symmetry called quasisymmetry (QS)[10, 11], forces the magnetic field strength, $B=|\mathbf{B}|$, to be symmetric [10, 12], whereas $\bm{B}$ itself depends on all three spatial coordinates, and is thus said to be three-dimensional (3D). The similarity of this symmetry with particle relabelling symmetries [13, 14] is known [12, 15] but yet to be fully explored. The hidden symmetry in $B$ makes the bounce action (second adiabatic invariant), $J_{\|}$, independent of the field-line label [11], which guarantees excellent particle confinement on a torus. The three-dimensional nature of B allows stellarators to avoid intrinsic difficulties (such as powerful transients) associated with large plasma currents in axisymmetric configurations.
It is currently unknown whether QS can be achieved exactly in a toroidal volume for a magnetic field that satisfies the constraint of force balance in ideal magnetohydrodynamics (MHD) [16]. The problem is of fundamental importance and potentially significant for the design of next-generation stellarator fusion reactors. Analytical results [17, 18, 19, 20] seem to indicate that in the presence of scalar plasma pressure, the resulting 3D nonlinear system of PDEs does not have solutions in a toroidal volume [17, 18, 20, 21, 22, 23]. In particular, the near-axis expansion (NAE), which is an asymptotic expansion in the distance from the magnetic axis [24, 25, 26], when applied to QS [17, 18, 27], becomes overdetermined beyond the second-order. Moreover, NAEs are typically divergent beyond second-order (L. Fu, private communication). Overdetermined problems do not generally have solutions. Even when they have solutions, they can be subject to restrictive constraints. For example, in the particular case of isodynamic magnetic fields, which satisfy a constraint more restrictive than QS [11, 28], it has been shown [29, 28] that if a solution exists, it must be essentially 2D: either axisymmetric like a tokamak or a helically symmetric straight stellarator.
@@
In this work, we demonstrate that QS admits a class of analytic $B$ that is a periodic solution [9, 30] of the KdV equation. We proceed along three independent lines of inquiry to support our result. In the first approach, we formulate the overdetermined problem of QS in a vacuum field in close analogy with the equations that arise in the theory of shallow water waves. Analogously to the water wave theory, we highlight the importance of a traveling wave (TW) frame and the existence of longer length scales in QS. Using a weakly nonlinear multiscale approach, we show that under some assumptions, the magnetic field strength must satisfy the KdV equation whose solutions satisfy the overdetermined set of equations perturbatively. We can also show that for small rotational transform, the next order corrections that are needed will lead to Gardner’s equation, a generalization of the KdV equation [31, 32, 33]. As in the isodynamic case [28], we show that the rotational transform is not free but highly constrained, and that for a given rotational transform, a perturbative solution of the overdetermination problem of quasisymmetric vacuum fields can be constructed only for some values of the shear. Although the perturbative approach is necessarily approximate, it provides information not only on quasisymmetric $B$ but also the full $\bm{B}$ over the entire plasma volume.
@@
Our third approach is data-driven. We use PySINDy[34, 35] to look for relations between $B$ and its derivatives in the large family of quasi-axisymmetric configurations generated in [36] and the QUASR database [37]. PySINDy employs sparse regression techniques and algorithms to discover governing dynamical systems models from noisy and limited data. We show that over a wide range of configurations and parameters, the KdV equation with the traveling wave form for $B$ holds. As described later in the paper, the only exceptions we found are covered by Gardner’s equation.
The connection with periodic KdV solitons allows us to describe $B$ on a magnetic flux surface with only three functions of the magnetic flux, related to the so-called spectral parameters [9] of the KdV equation. We show that these functions are closely related to the spectrum of $B$ in Boozer coordinates [11] where the hidden symmetry is manifest, thereby highlighting the low dimensionality of the Boozer spectrum. Furthermore, our analysis shows that the periodicity of $B$ is maintained only when these functions are distinct. Consequently, we can estimate the region of validity of volumetric QS based on where the two spectral parameters intersect (discussed below). Remarkably, near the point of intersection, the connection length (period of $B$) begins to diverge, and $B$ approaches the form of a reflectionless potential (a single soliton that decays at infinity) [38, 39]. We demonstrate using both data-driven and traditional numerical methods that our analytic theory can describe well the properties of the field strength on a flux surface of the vacuum equilibrium with precise quasi-axisymmetry (QA)[4] and many other optimized quasisymmetric configurations.
@@ II Basic formulation of quasisymmetry
We begin by reviewing some of the known basic properties of QS. Several equivalent [11, 12, 16, 40, 22] mathematical descriptions of QS exist. In this Section, we will briefly discuss the most relevant equations and coordinates to describe QS and point out their inter-relationships.
It can be shown that QS implies the existence of a divergence-free vector field $\bm{u}$ that satisfies the overdetermined system [22]
@@
The mathematical structure of (1) strongly suggests that $\bm{u}$ can be associated with a particle relabelling symmetry group ([13, 41, 42, 43, 44, 14] and references therein) with density replaced by the field strength [12, 15]. Therefore, understanding the hidden symmetry behind QS might also be beneficial in understanding Casimirs [45, 13] in fluid dynamics and ideal MHD.
@@
which is sometimes called the two-term form [22].
For a vacuum field $\bm{B}=G_{0}\bm{\nabla}\Phi$ [46, 11],
@@
where $\iota$ is the rotational transform, $N$ is the helicity of QS, $2\pi G_{0}/\mu_{0}$ is the poloidal current outside the surface and $\Phi$ is a toroidal angle. Moreover, the two-term form of QS (equivalently $\bm{u}\cdot\bm{\nabla}B=0$) for vacuum fields leads to a TW solution for $B$ [47] of the form
@@
Another form of QS is the so-called “triple-product” form [12, 22]
@@
The two-term and the triple-product form of QS are equivalent under the assumption of irrational $\iota$ and continuity on rational surfaces [11]. Alternatively, the form
@@
It is useful to use the standard Clebsch coordinate system [11, 48] , $(\psi,\alpha,\ell)$, such that $\bm{B}=\bm{\nabla}\psi\times\bm{\nabla}\alpha$ with flux label $\psi$, field-line label $\alpha$, and $\ell$, the arclength along the magnetic field. We impose suitable toroidal cuts since $\alpha$ is a multi-valued function. The triple-product form (8) now reads [16]
@@
where $f$ must be a function of $\psi$ and $B$.
In toroidal geometry, $B$ is subject to the following periodic boundary condition [11]:
@@
The period of $B$, $L(\psi)$, is called the connection length and is a quantity of great physical interest [49].
As shown in Appendix A, in $(\psi,\alpha,\ell)$ coordinates, the $\bm{u}\cdot\bm{\nabla}B=0$ condition takes the form [16]
@@
Thus, there exists a special TW frame (¯$\overline{\ell},\alpha$) in QS where the field strength $B$ is independent of $\alpha$ [16]. We show the equivalence of the two TW forms of $B$ given by (7) and (16) in Appendix B.1.
@@
where $\mathcal{E}$ is related to the particle energy, is a special case of (17). As a direct consequence of (17), it follows that quasisymmetric $B$ can not have local maxima or minima on a flux surface [11, 50].
The distinctions between the cases of exact symmetry (axisymmetry or helical symmetry), quasisymmetry and omnigeneity must now be made clear. For the special cases of exact symmetry, the corresponding symmetry vectors are the well-known Killing vectors of Euclidean space [40], which implies that $\bm{u}$ and $\bm{B}$ are completely decoupled, and the symmetric form of $\bm{B}$ is determined by $\bm{u}$. Therefore, the condition (1b) can be trivially satisfied. Here, we will focus on the case of 3D QS where $\bm{u}$ cannot be posited a priori. Consequently, the condition (1b) will impose a nontrivial constraint. In particular, unlike the case of exact symmetry, the functional form of $f(B,\psi)$ in the $\partial_{\ell}B$ equation will be constrained. We address this question directly in this work.
QS being a special case of omnigeneity [11], the integral property (17) also holds for omnigeneous magnetic fields. However, significant differences arise because omnigeneity is not a local property, and a symmetry vector $\bm{u}$ does not exist for omnigeneous magnetic fields. Thus, the local conditions such as (4) or (10), which follow from the properties of $\bm{u}$ do not hold for omnigeneity. It follows that omnigenous fields need not be of the TW form (6). Furthermore, unlike QS [17, 27], omnigeneous $B$ cannot be uniformly analytic on a flux surface [51, 52, 53].
As a final step, one must impose a force-balance condition. Without imposing force balance, one can exactly satisfy the QS condition order by order in the NAE [23]. However, the nonlinear overdetermination problem that results when the force-balance condition is imposed is severely restrictive [17, 18]. In the next Section, we address this problem head-on by looking at quasisymmetric vacuum fields. To make the problem analytically tractable, we first consider slab geometry (or a flat torus). We then show that the analysis for the general geometry can be carried out in close analogy with the slab geometry under certain additional assumptions.
III Perturbative solution to the overdetermination problem of QS in vacuum magnetic fields

### v3

@@ Abstract
Quasisymmetry (QS) is a hidden symmetry of the magnetic field strength, $|\mathbf{B}|$, that confines charged particles effectively in a three-dimensional toroidal plasma equilibrium. Here, we show that QS has a deep connection to the underlying symmetry that makes solitons possible. Our approach uncovers a hidden lower dimensionality of $|\mathbf{B}|$ on a magnetic flux surface, which could make stellarator optimization schemes significantly more efficient. Recent numerical breakthroughs (M. Landreman and E. Paul, Phys. Rev. Lett. 128, 035001 (2022)) have yielded configurations with excellent volumetric QS and surprisingly low magnetic shear. Our approach elucidates why the magnetic shear is low in these configurations. Furthermore, given $|\mathbf{B}|$, we deduce an upper bound on the maximum toroidal volume that can be quasisymmetric only from the properties of $|\mathbf{B}|$ and verify it for the Landreman-Paul precise quasiaxisymmetric (QA) stellarator configuration. In the neighborhood of the outermost surface, we show that $|\mathbf{B}|$ approaches the form of the 1-soliton reflectionless potential (I. Gjaja and A. Bhattacharjee, Phys. Rev. Lett. 68, 2413 (1992)). The connection length diverges, indicating the possible presence of an X-point that could potentially be used as basis for a divertor. We present three independent approaches to demonstrate that quasisymmetric $|\mathbf{B}|$ is described by well-known integrable systems such as the Korteweg-de Vries (KdV) equation. The first approach is weakly nonlinear multiscale perturbation theory, which highlights the crucial role that magnetic shear plays in QS. We show that the overdetermined problem of finding quasisymmetric vacuum fields admits solutions for which the rotational transform is highly constrained. We obtain the KdV equation (and, more specifically, Gardner’s equation for certain choices of parameters). Our second approach is non-perturbative and based on ensuring single-valuedness of $|\bm{B}|$, which directly leads to its Painlevé property shared by the KdV equation. Our third approach uses machine learning, trained on a large dataset of numerically optimized quasisymmetric stellarators. We robustly recover the KdV (and Gardner’s) equation from the data.
@@ I Introduction
Physical systems possessing a hidden symmetry attract serious attention Charalambous et al. (2021); Lee et al. (2021); Liu and Tegmark (2022); Landreman and Paul (2022); Moser (1979); Leach and Andriopoulos (2008) because of the far-reaching impact of the symmetry on the physical properties of the systems. The hidden symmetry manifests only in special coordinates, which are constructed, along with the invariants, as part of the solution to the problem. In the presence of hidden symmetries, the partial differential equations (PDEs) governing classical dynamical fields often reduce to integrable PDEs (Moser, 1979; Ovsienko and Khesin, 1987; Leach and Andriopoulos, 2008). An example of the latter is the well-known Korteweg-de Vries (KdV) equation with soliton solutions Gardner et al. (1967); Novikov et al. (1984).
In plasma physics, a hidden symmetry called quasisymmetry (QS)(Boozer, 1983; Helander, 2014), forces the magnetic field strength, $B=|\mathbf{B}|$, to be symmetric (Boozer, 1983; Rodriguez, 2022), whereas $\bm{B}$ itself depends on all three spatial coordinates, and is thus said to be three-dimensional (3D). The similarity of this symmetry with particle relabelling symmetries (Padhye and Morrison, 1996a; Webb, 2018) is known (Rodriguez, 2022; Weitzner and Sengupta, 2021) but yet to be fully explored. The hidden symmetry in $B$ makes the bounce action (second adiabatic invariant), $J_{\|}$, independent of the field-line label (Helander, 2014), which guarantees excellent particle confinement in a torus. The 3D nature of B allows stellarators to avoid intrinsic difficulties (such as powerful transients) associated with large plasma currents in axisymmetric configurations such as tokamaks.
It is currently unknown whether exact QS can be achieved exactly in a toroidal volume for a magnetic field that satisfies the constraint of force balance in ideal magnetohydrodynamics (MHD) Freidberg (2014). The problem is of fundamental importance and potentially significant for the design of next-generation stellarator fusion reactors. Analytical results (Garren and Boozer, 1991a, b; Plunk and Helander, 2018; Landreman and Sengupta, 2019) seem to indicate that in the presence of scalar plasma pressure, the resulting 3D nonlinear system of PDEs does not have solutions in a toroidal volume (Garren and Boozer, 1991a, b; Landreman and Sengupta, 2019; Constantin et al., 2021; Rodríguez et al., 2020, 2022). In particular, the near-axis expansion (NAE), which is an asymptotic expansion in the distance from the magnetic axis (Mercier, 1974; Solov’ev and Shafranov, 1970; Bernardin et al., 1986), when applied to QS (Garren and Boozer, 1991a, b; Landreman and Sengupta, 2018), becomes overdetermined beyond the second-order. Moreover, NAEs are typically divergent beyond second-order (L. Fu, private communication). Overdetermined problems do not generally have solutions. Even when they have solutions, they can be subject to restrictive constraints. For example, in the particular case of isodynamic magnetic fields, which satisfy a constraint more restrictive than QS (Helander, 2014; Schief, 2003), it has been shown (Palumbo, 1968; Schief, 2003) that if a solution exists, it must be essentially 2D: either axisymmetric like a tokamak or a helically symmetric straight stellarator.
@@
In this work, we demonstrate that QS admits a class of analytic $B$ that is a periodic solution (Novikov et al., 1984; Kamchatnov, 2000) of the KdV equation. We proceed along three independent lines of inquiry to support our result. In the first approach, we formulate the overdetermined problem of QS in a vacuum field in close analogy with the equations that arise in the theory of shallow water waves. Analogously to the water wave theory, we highlight the importance of a traveling wave (TW) frame and the existence of longer length scales in QS. Using a weakly nonlinear multiscale perturbation theory, we show that under some assumptions, the magnetic field strength must satisfy the KdV equation whose solutions satisfy the overdetermined set of equations. We can also show that for small rotational transform, the next order corrections that are needed will lead to Gardner’s equation, a generalization of the KdV equation (Miura et al., 1968; Su and Gardner, 1969; Kamchatnov et al., 2012). As in the isodynamic case (Schief, 2003), we show that the rotational transform is not free but highly constrained, and that for a given rotational transform, a perturbative solution of the overdetermination problem of quasisymmetric vacuum fields can be constructed only for some values of the shear. Although the perturbative approach is necessarily approximate, it provides information not only on quasisymmetric $B$ but also the full $\bm{B}$ over the entire plasma volume.
@@
Our third approach is data-driven. We use PySINDyde Silva et al. (2020); Kaptanoglu et al. (2022) to look for relations between $B$ and its derivatives in the large family of quasi-axisymmetric configurations generated in Buller et al. (2024) and the QUASR database Giuliani (2023). PySINDy employs sparse regression techniques and algorithms to discover governing dynamical systems models from noisy and limited data. We show that over a wide range of configurations and parameters, the KdV equation with the traveling wave form for $B$ holds. As described later in the paper, the only exceptions we found are covered by Gardner’s equation.
The connection with periodic KdV solitons allows us to describe $B$ on a magnetic flux surface with only three functions of the magnetic flux, related to the so-called spectral parameters (Novikov et al., 1984) of the KdV equation. We show that these functions are closely related to the spectrum of $B$ in Boozer coordinates Helander (2014) where the hidden symmetry is manifest, thereby highlighting the low dimensionality of the Boozer spectrum. Furthermore, our analysis shows that the periodicity of $B$ is maintained only when these functions are distinct. Consequently, we can estimate the region of validity of volumetric QS based on where the two spectral parameters intersect (discussed below). Remarkably, near the point of intersection, the connection length (period of $B$) begins to diverge, and $B$ approaches the form of a reflectionless potential (a single soliton that decays at infinity) (Gjaja and Bhattacharjee, 1992; Berry and Howls, 1990). The diverging connection length is a characteristic of an X-point that could potentially be the basis for a non-resonant divertor (Mioduszewski et al., 2007; Boozer, 2015; Punjabi and Boozer, 2020).
We demonstrate using both data-driven and traditional numerical methods that our analytic theory can describe well the properties of the field strength on a flux surface of the vacuum equilibrium with precise quasi-axisymmetry (QA)Landreman and Paul (2022) and many other optimized quasisymmetric configurations.
@@ II Basic formulation of quasisymmetry
We begin by reviewing some of the known basic properties of QS. Several equivalent (Helander, 2014; Rodriguez, 2022; Freidberg, 2014; Burby et al., 2020; Rodríguez et al., 2020) mathematical descriptions of QS exist. In this Section, we will briefly discuss the most relevant equations and coordinates to describe QS and point out their inter-relationships.
It can be shown that QS implies the existence of a divergence-free vector field $\bm{u}$ that satisfies the system (Rodríguez et al., 2020)
@@
The mathematical structure of (1) strongly suggests that $\bm{u}$ can be associated with a particle relabelling symmetry group (Padhye and Morrison (1996a, b); Salmon (1982, 1988); Webb and Zank (2006); Webb (2018) and references therein) with density replaced by the field strength (Rodriguez, 2022; Weitzner and Sengupta, 2021). Therefore, understanding the hidden symmetry underlying QS might also be beneficial in understanding Casimirs (Hameiri, 2004; Padhye and Morrison, 1996a) in fluid dynamics and ideal MHD.
@@
which is sometimes called the two-term form of QS Rodríguez et al. (2020).
For a vacuum field $\bm{B}=G_{0}\bm{\nabla}\Phi$ Boozer (1980); Helander (2014),
@@
where $\iota$ is the rotational transform, $N$ is the helicity of QS, $2\pi G_{0}/\mu_{0}$ is the poloidal current outside the surface and $\Phi$ is a toroidal angle. Moreover, the two-term form of QS (equivalently $\bm{u}\cdot\bm{\nabla}B=0$) for vacuum fields leads to a TW solution for $B$ (Sengupta et al., 2021) of the form
@@
Another form of QS is the so-called “triple-product” form (Rodriguez, 2022; Rodríguez et al., 2020)
@@
The two-term and the triple-product form of QS are equivalent under the assumption of irrational $\iota$ and continuity on rational surfaces (Helander, 2014). Alternatively, the form
@@
It is useful to use the standard Clebsch coordinate system Helander (2014); D’haeseleer et al. (2012) , $(\psi,\alpha,\ell)$, such that $\bm{B}=\bm{\nabla}\psi\times\bm{\nabla}\alpha$ with flux label $\psi$, field-line label $\alpha$, and $\ell$, the arclength along the magnetic field. We impose suitable toroidal cuts since $\alpha$ is a multi-valued function. The triple-product form (8) now reads (Freidberg, 2014)
@@
where $f$ must be a function of $\psi$ and $B$. Note that because $\partial_{\ell}B$ can be both positive and negative, we will assume that $f(B,\psi)$ also depends on $\sigma=\text{sign}(\partial_{\ell}B)$.
In toroidal geometry, $B$ is subject to the following periodic boundary condition (Helander, 2014):
@@
The period of $B$, $L(\psi)$, is called the connection length and is a quantity of great physical interest Barnes et al. (2011).
As shown in Appendix A, in $(\psi,\alpha,\ell)$ coordinates, the $\bm{u}\cdot\bm{\nabla}B=0$ condition takes the form Freidberg (2014)
@@
Thus, there exists a special TW frame (¯$\overline{\ell},\alpha$) in QS where the field strength $B$ is independent of $\alpha$ (Freidberg, 2014). We show the equivalence of the two TW forms of $B$ given by (7) and (16) in Appendix B.1.
@@
where $\mathcal{E}$ is related to the particle energy, is a special case of (17). It follows from (17) that quasisymmetric $B$ can not have local maxima or minima on a flux surface (Helander, 2014; Landreman and Catto, 2012).
The distinctions between the cases of exact symmetry (axisymmetry or helical symmetry), quasisymmetry and omnigeneity must now be made clear. For the special cases of exact symmetry, the corresponding symmetry vectors are the well-known Killing vectors of Euclidean space (Burby et al., 2020), which implies that $\bm{u}$ and $\bm{B}$ are completely decoupled, and the symmetric form of $\bm{B}$ is determined by $\bm{u}$. Therefore, the condition (1b) can be trivially satisfied. Here, we will focus on the case of 3D QS where $\bm{u}$ cannot be posited a priori. Consequently, the condition (1b) will impose a nontrivial constraint. In particular, unlike the case of exact symmetry, the functional form of $f(B,\psi)$ in the $\partial_{\ell}B$ equation must be constrained. We address this question directly in this work.
QS being a special case of omnigeneity (Helander, 2014), the integral property (17) also holds for omnigeneous magnetic fields. However, significant differences arise because omnigeneity is not a local property, and a symmetry vector $\bm{u}$ does not exist for omnigeneous magnetic fields. Thus, the local conditions such as (4) or (10), which follow from the properties of $\bm{u}$ do not hold for omnigeneity. It follows that omnigenous fields need not be of the TW form (6). Furthermore, unlike QS (Garren and Boozer, 1991a; Landreman and Sengupta, 2018), omnigeneous $B$ cannot be uniformly analytic on a flux surface (Cary and Shasharina, 1997a, b; Plunk et al., 2019).
As a final step, one must impose a force-balance condition. Without imposing force balance, one can satisfy the QS condition exactly order by order in the NAE (Rodríguez et al., 2022). However, the nonlinear overdetermination problem that results when the force-balance condition is imposed is severely restrictive (Garren and Boozer, 1991a, b). In the next Section, we address this problem head-on by looking at quasisymmetric vacuum fields. To make the problem analytically tractable, we first consider slab geometry (or a flat torus). We then show that the analysis for the general geometry can be carried out in close analogy with the slab geometry under certain additional assumptions.
III Perturbative solution to the overdetermination problem of QS for vacuum magnetic fields